OpenStudy (anonymous):

write -2/3x +1/2y = 1 in the general form with integer coefficients.. 4x - 3y +6 =0 is my answer

OpenStudy (anonymous):

-2/3x +1/2y = 1 ===> -4x+3y =6 ===> 4x-3y+6=0 ===> your answer is correct

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you are welcome.